# From Prompt Injections to Protocol Exploits: Threats in LLM-Powered AI Agents Workflows

### v1

@@ II-B Security Challenges and Threat Models
Deng et al. [9] present a focused survey on the emerging security threats associated with artificial intelligence (AI) agents, which are autonomous software systems capable of perceiving inputs, planning tasks, and executing actions. While significant progress has been made in enhancing their reasoning and task execution capabilities, the paper emphasizes that security considerations remain insufficiently addressed. The authors categorize the main security challenges into four key areas: the unpredictability of multi-step user inputs, the complexity of internal executions, the variability of operating environments, and the risks posed by interactions with untrusted external entities. Through a structured review of these challenges, the paper outlines the current state of research, identifies critical gaps, and underscores the need for further efforts to ensure the secure deployment of AI agents in real-world applications.
Wang et al. [27] present an in-depth and pioneering survey that introduces the concept of “full-stack” safety for Large Language Models (LLMs), addressing security and safety concerns across the entire lifecycle of these models. While most existing studies focus narrowly on specific phases such as deployment or fine-tuning, this work offers a holistic perspective that spans data preparation, pre-training, post-training, deployment, and commercialization. Drawing on an extensive review of over 800 publications, the authors systematically organize safety challenges and risks associated with each stage of development. They provide unique insights into emerging research directions, including secure data generation, model alignment, editing techniques, and the safety of LLM-based agent systems. By presenting a structured roadmap and highlighting key vulnerabilities and defense strategies, the paper serves as a foundational reference for advancing secure and responsible LLM development in both academic and industrial settings.
@@ II-C LLMs in Multi-Agent and Interactive Systems
Yan et al. [5] offer a comprehensive survey that explores the integration of Large Language Models (LLMs) into multi-agent systems (MAS), emphasizing the central role of communication in enabling collaborative and intelligent behavior. By analyzing system-level features, such as architectural design and communication objectives, along with internal mechanisms including strategies, paradigms, and message content, the paper highlights how LLM-based agents coordinate and exchange information to achieve complex tasks that exceed the capabilities of individual agents. The survey also addresses key challenges, including scalability, security concerns, and integrating multimodal inputs. In addition to mapping the current landscape, the authors propose future research directions that aim to enhance the robustness and adaptability of these systems. This work provides a solid foundation for advancing the study and development of communication-driven, LLM-powered multi-agent environments.
Chen et al. [28] explore the growing landscape of Computer-Using Agents (CUAs), a new class of AI-driven systems that autonomously interact with graphical user interfaces to perform tasks across desktop, web, and mobile environments. As these agents evolve from simple prototypes to complex LLM-based systems, they bring an expanded surface for safety and security vulnerabilities. The paper highlights challenges introduced by integrating multimodal inputs and the reasoning limitations of large language models. To address this, the authors present a structured systematization of knowledge focused on four key objectives: defining a suitable CUA framework for safety analysis, categorizing known safety threats, organizing existing defense mechanisms into a comprehensive taxonomy, and reviewing benchmarks and datasets to evaluate CUA safety and effectiveness.
@@ II-D Applications and Benchmarking of LLM Agents
Ferrag et al. [2] present a comprehensive and much-needed synthesis of the rapidly evolving ecosystem surrounding large language models and autonomous AI agents. Recognizing the fragmented nature of existing evaluation benchmarks and agent frameworks, they provide a detailed side-by-side comparison of benchmarks developed between 2019 and 2025, spanning diverse domains such as general reasoning, code generation, factual retrieval, and multimodal tasks. The paper introduces a structured taxonomy comprising approximately 60 benchmarks, offering clarity across various areas, including software engineering, task orchestration, and interactive assessment. Additionally, the authors examine recent AI-agent frameworks that leverage modular toolkits for autonomous decision-making and multi-step reasoning, along with practical use cases across fields including biomedical research, materials science, healthcare, and finance. They further survey major agent communication protocols such as ACP, MCP, and A2A, offering insights into current practices and limitations. Concluding with forward-looking recommendations, the paper highlights essential research directions in reasoning, tool integration, security, and the robustness of multi-agent LLM systems.
@@ II-E Edge Intelligence and LLM Integration
Friha et al. [29] present a thorough and forward-looking survey on the integration of Large Language Models (LLMs) with Edge Intelligence (EI), framing it as a transformative approach for enabling advanced capabilities on decentralized edge devices. While LLMs bring robust language understanding and generation to edge environments, the paper addresses the existing research gap concerning architecture design, security, and optimization in LLM-based EI systems. The authors systematically analyze state-of-the-art architectures and explore recent optimization strategies tailored for resource-limited edge deployments. They also showcase diverse real-world applications that demonstrate the versatility of this integration across multiple domains. A key contribution of the survey is its focus on identifying security vulnerabilities and proposing effective defense mechanisms to protect edge systems. Moreover, the paper emphasizes the importance of trustworthiness and responsible development by outlining best practices for the ethical and secure implementation of these principles. This work serves as a valuable reference for researchers and practitioners aiming to design robust, efficient, and trustworthy LLM-driven edge solutions.
@@ II-F LLMs for Cybersecurity Applications
Ferrag et al. [30] present a comprehensive survey on the evolving role of Generative AI and large language models (LLMs) in shaping the future of cybersecurity. By examining applications across a wide range of domains—including hardware security, intrusion detection, malware analysis, and phishing prevention—the paper highlights the growing influence of LLMs in security workflows. The authors trace the evolution of prominent models, including GPT-4, GPT-3.5, Mixtral-8x7B, Falcon2, BERT, and LLaMA, while also identifying critical vulnerabilities such as prompt injection, data poisoning, and adversarial misuse. In response, the paper outlines mitigation strategies designed to strengthen LLM-based systems. A notable contribution is the performance evaluation of 42 LLMs in cybersecurity-specific tasks, alongside a detailed review of available datasets and their limitations. Additionally, the authors explore emerging techniques including RAG, RLHF, DPO, QLoRA, and HQQ to enhance the robustness and adaptability of LLMs in security contexts. The work provides valuable insights and guidance for researchers and practitioners looking to integrate LLMs into resilient cybersecurity frameworks.
@@ II-G Comparison with Existing Surveys
While previous surveys have offered valuable insights into communication protocols [13, 25], agent architectures [28], and broad security concerns in LLM-based systems [9, 27, 29, 30], none have comprehensively examined the specific vulnerabilities and attack techniques that emerge within LLM-powered AI agent communications. Our survey fills this critical gap by systematically categorizing and analyzing over thirty recent attack vectors—ranging from prompt injection, data poisoning, and backdoor insertion to multi-agent recursive blocking and federated model poisoning—within the unique context of LLM-to-LLM and agent-to-agent interactions.
Unlike prior works, which often address either generic LLM safety [27], tool integration [2], or security in isolated modalities [31, 32], our survey provides a vertical deep dive into vulnerabilities across multiple communication layers (e.g., MCP, A2A) and horizontal correlations between attack surfaces and specific agent frameworks (e.g., CrewAI, LangChain, AutoGen). Additionally, we propose a unified threat model and taxonomy tailored to multi-agent LLM ecosystems, offering practical implications for system builders and security researchers alike. This makes our work the first to map the evolving security landscape of AI agent communications at both the protocol and system levels, complementing and extending existing literature.
@@
Perez et al.[33] 2022 Prompt‐Based Direct Prompt Injection (PromptInject) GPT‐3 (transformer‐based LLM) Discussion of need for stronger protective measures; no concrete defense tested Simple handcrafted inputs can induce goal hijacking and prompt leaking; low‐skill attackers can exploit stochastic behavior to misalign GPT‐3
@@
Pedro et al.[34] 2023 Prompt‐Based Prompt‐to‐SQL (P2SQL) Injection Langchain‐based web apps integrating LLMs Four defense strategies integrated into Langchain (e.g., input sanitization, query validation) Seven SOTA LLMs are widely susceptible; defense strategies significantly reduce the success of SQL injection via prompts
@@
Bagdasaryan et al.[35] 2023 Prompt‐Based Indirect & Compositional Prompt Injection Multimodal LLMs (LLaVA, PandaGPT) Recommendation for enhanced detection of adversarial perturbations in images/audio Adversarial perturbations embedded in images or sounds can steer multimodal LLM outputs without modifying textual prompt
@@
Zhan et al.[16] 2025 Prompt‐Based Adaptive Indirect Prompt Injection LLM agents using external tools Emphasis on designing defenses resilient to adaptive attacks; no novel defense, but evaluation methodology Eight existing defenses bypassed with adaptive strategies; attack success rate >50% for each defense
@@
Jiang et al.[17] 2025 Jailbreaking Compositional Instruction Attack (CIA) GPT‐4, GPT‐3.5, Llama2‐70B Intent‐Based Defense (IBD) to detect harmful intents CIA achieves >95% ASR; IBD reduces ASR by >74% by identifying underlying malicious intent
@@
Wei et al.[36] 2023 Jailbreaking In‐Context Demonstrations Attack (ICA) Various aligned LLMs In‐Context Defense (ICD) providing refusal examples Harmful demonstrations can increase jailbreak success; ICD effectively counters ICA by reducing success rates
@@
Shayegani et al.[37] 2023 Jailbreaking Cross‐Modality Jailbreak Attack Vision‐Language Models (e.g., CLIP‐integrated VLMs) Need for new alignment strategies in multi‐modal systems Adversarial images targeting toxic embeddings, combined with text, can break VLM alignment without model access
@@
Anil et al.[38] 2024 Jailbreaking Long‐Context Jailbreak Attack Closed‐weight LLMs (e.g., Anthropic, OpenAI) Recommendation for enhanced safety in long‐context models Attack success follows a power law as demonstration count increases; hundreds of shots suffice to manipulate behavior
@@
Liu et al.[39] 2023 Jailbreaking Automated Jailbreak Prompt Generation (AutoDAN) Aligned LLMs fine‐tuned with human feedback Hierarchical genetic algorithm framework AutoDAN produces stealthy, semantically meaningful prompts; outperforms prior methods in cross‐model transfer and stealth
@@
Yu et al.[18] 2023 Jailbreaking Jailbreak Fuzzing (GPTFuzz) ChatGPT, LLaMA‐2, Vicuna, etc. Judgment model to evaluate and refine templates >90% success rates across multiple LLMs; framework scales better than manual template crafting
@@
Schwartz et al.[19] 2025 Jailbreaking Graph of Attacks with Pruning (GAP) GPT‐3.5 Automatic seed prompt generation; PromptGuard fine‐tuning for detection GAP achieves 92% ASR using 54% of queries required before; PromptGuard accuracy improves from 5.1% to 93.89% on Toxic‐Chat
@@
Dong et al.[40] 2025 Jailbreaking JailFuzzer (LLM‐Agent‐Driven) Text‐to‐Image generative models LLM‐based guided mutation and oracle evaluation High success rates with minimal queries; generates natural, semantically coherent jailbreak prompts in black‐box settings
@@
Wang et al.[41] 2023 Adversarial & Evasion adversarial In‐Context Learning (advICL) In‐Context Learning in LLMs Transferable‐advICL to craft demos for unseen inputs Robustness decreases as demonstration count increases; advICL and Transferable‐advICL mislead model on unseen inputs
@@
Zhuang et al.[42] 2023 Adversarial & Evasion Query‐Free Adversarial Attack Stable Diffusion (T2I model with CLIP encoder) Insights on text encoder robustness; suggests encoder hardening Five‐character perturbations can drastically alter generated images; targeted attacks steer specific content manipulations
@@
Yang et al.[31] 2024 Adversarial & Evasion Multimodal Adversarial Attack (MMA‐Diffusion) T2I Models (open and commercial) Need for combined textual and visual defense strategies Bypasses prompt filters and post‐hoc safety checkers using text + visual manipulation; NSFW content generated despite safeguards
@@
Chen et al.[20] 2025 Adversarial & Evasion Active Environment Injection Attack (AEIA) MLLM‐Based Agents in Mobile OS AEIA‐MN framework to evaluate and reveal vulnerabilities Maximum 93% ASR on AndroidWorld when combining interface and reasoning gaps; advanced MLLMs remain highly vulnerable
@@
Patlan et al.[43] 2025 Adversarial & Evasion Context Manipulation Attack Web3‐Based AI Agents Fine‐tuning methods more robust than prompt defenses CrAIBench finds prompt defenses ineffective; fine‐tuning reduces ASR significantly while preserving utility in single‐step tasks
@@
Input Manipulation Attacks Prompt-Based Attacks Direct Prompt Injection (Perez et al. [33]) Prompt-to-SQL Injection (Pedro et al. [34]) Indirect & Compositional Prompt Injection (Bagdasaryan et al. [35]) Adaptive Indirect Prompt Injection (Zhan et al. [16]) Toxic Agent Flow Attack (Invariant Labs [invariantlabs2025mcp]) Jailbreaking Attacks Compositional Instruction Attack (Jiang et al. [17]) In-Context Demonstrations (Wei et al. [36]) Long-Context Attacks (Anil et al. [38]) Automated Jailbreak Prompt Generation (Liu et al. [39]) Jailbreak-Fuzzing Attack (Dong et al. [40]) Adversarial & Evasion Attacks Adv-ICL Attacks (Wang et al. [41]) Query-Free Attack (Zhuang et al. [42]) Multimodal Adversarial Attack (Yang et al. [31]) Active Environment Injection Attack (Chen et al. [20]) Context Manipulation Attack (Patlan et al. [43])

### v2

@@ II-B Security Challenges and Threat Models
Deng et al. [deng2025ai] present a focused survey on the emerging security threats associated with artificial intelligence (AI) agents, which are autonomous software systems capable of perceiving inputs, planning tasks, and executing actions. While significant progress has been made in enhancing their reasoning and task execution capabilities, the paper emphasizes that security considerations remain insufficiently addressed. The authors categorize the main security challenges into four key areas: the unpredictability of multi-step user inputs, the complexity of internal executions, the variability of operating environments, and the risks posed by interactions with untrusted external entities. Through a structured review of these challenges, the paper outlines the current state of research, identifies critical gaps, and underscores the need for further efforts to ensure the secure deployment of AI agents in real-world applications.
Wang et al. [wang2025comprehensive] present an in-depth, pioneering survey that introduces the concept of “full-stack” safety for Large Language Models (LLMs), addressing security and safety concerns throughout their lifecycle. While most existing studies focus narrowly on specific phases, such as deployment or fine-tuning, this work offers a holistic perspective spanning data preparation, pre-training, post-training, deployment, and commercialization. Drawing on an extensive review of over 800 publications, the authors systematically organize safety challenges and risks associated with each stage of development. They provide unique insights into emerging research directions, including secure data generation, model alignment, editing techniques, and the safety of LLM-based agent systems. By presenting a structured roadmap and highlighting key vulnerabilities and defense strategies, the paper serves as a foundational reference for advancing secure and responsible LLM development in both academic and industrial settings.
@@ II-C LLMs in Multi-Agent and Interactive Systems
Yan et al. [yan2025beyond] offer a comprehensive survey that explores the integration of Large Language Models (LLMs) into multi-agent systems (MAS), emphasizing the central role of communication in enabling collaborative and intelligent behavior. By analyzing system-level features —such as architectural design and communication objectives —and internal mechanisms, including strategies, paradigms, and message content, the paper highlights how LLM-based agents coordinate and exchange information to achieve complex tasks that exceed the capabilities of individual agents. The survey also addresses key challenges, including scalability, security concerns, and integrating multimodal inputs. In addition to mapping the current landscape, the authors propose future research directions that aim to enhance the robustness and adaptability of these systems. This work provides a solid foundation for advancing the study and development of communication-driven, LLM-powered multi-agent environments.
Chen et al. [chen2025survey] explore the growing landscape of Computer-Using Agents (CUAs), a new class of AI-driven systems that autonomously interact with graphical user interfaces to perform tasks across desktop, web, and mobile environments. As these agents evolve from simple prototypes to complex LLM-based systems, they create a larger surface area for safety and security vulnerabilities. The paper highlights challenges introduced by integrating multimodal inputs and the limitations of large language models in reasoning. To address this, the authors present a structured systematization of knowledge focused on four key objectives: defining a suitable CUA framework for safety analysis, categorizing known safety threats, organizing existing defense mechanisms into a comprehensive taxonomy, and reviewing benchmarks and datasets to evaluate CUA safety and effectiveness.
@@ II-D Applications and Benchmarking of LLM Agents
Ferrag et al. [ferrag2025llm] present a comprehensive and much-needed synthesis of the rapidly evolving ecosystem surrounding large language models and autonomous AI agents. Recognizing the fragmented nature of existing evaluation benchmarks and agent frameworks, they provide a detailed side-by-side comparison of benchmarks developed between 2019 and 2025, spanning diverse domains such as general reasoning, code generation, factual retrieval, and multimodal tasks. The paper introduces a structured taxonomy comprising approximately 60 benchmarks that offers clarity across various areas, including software engineering, task orchestration, and interactive assessment. Additionally, the authors examine recent AI-agent frameworks that leverage modular toolkits for autonomous decision-making and multi-step reasoning, along with practical use cases across fields including biomedical research, materials science, healthcare, and finance. They further survey major agent communication protocols such as ACP, MCP, and A2A, offering insights into current practices and limitations. Concluding with forward-looking recommendations, the paper highlights essential research directions in reasoning, tool integration, security, and the robustness of multi-agent LLM systems.
@@ II-E Edge Intelligence and LLM Integration
Friha et al. [friha2024llm] present a thorough and forward-looking survey on the integration of Large Language Models (LLMs) with Edge Intelligence (EI), framing it as a transformative approach for enabling advanced capabilities on decentralized edge devices. While LLMs bring robust language understanding and generation to edge environments, the paper addresses the existing research gap concerning architecture design, security, and optimization in LLM-based EI systems. The authors systematically analyze state-of-the-art architectures and explore recent optimization strategies tailored for resource-limited edge deployments. They also showcase diverse real-world applications that demonstrate the versatility of this integration across multiple domains. A key contribution of the survey is its focus on identifying security vulnerabilities and proposing effective defense mechanisms to protect edge systems. Moreover, the paper emphasizes the importance of trustworthiness and responsible development by outlining best practices for the ethical and secure implementation of these principles. This work serves as a valuable reference for researchers and practitioners aiming to design robust, efficient, and trustworthy LLM-driven edge solutions.
@@ II-F LLMs for Cybersecurity Applications
Ferrag et al. [ferrag2025generative] present a comprehensive survey on the evolving role of Generative AI and large language models (LLMs) in shaping the future of cybersecurity. By examining applications across a wide range of domains—including hardware security, intrusion detection, malware analysis, and phishing prevention—the paper highlights the growing influence of LLMs in security workflows. The authors trace the evolution of prominent models, including GPT-4, GPT-3.5, Mixtral-8x7B, Falcon2, BERT, and LLaMA, and identify critical vulnerabilities such as prompt injection, data poisoning, and adversarial misuse. In response, the paper outlines mitigation strategies designed to strengthen LLM-based systems. A notable contribution is the performance evaluation of 42 LLMs across cybersecurity-specific tasks, along with a detailed review of available datasets and their limitations. Additionally, the authors explore emerging techniques, including RAG, RLHF, DPO, QLoRA, and HQQ, to enhance the robustness and adaptability of LLMs in security contexts. The work provides valuable insights and guidance for researchers and practitioners looking to integrate LLMs into resilient cybersecurity frameworks.
@@ II-G Comparison with Existing Surveys
While previous surveys have offered valuable insights into communication protocols [yang2025survey, hou2025model], agent architectures [chen2025survey], and broad security concerns in LLM-based systems [deng2025ai, wang2025comprehensive, friha2024llm, ferrag2025generative], none have comprehensively examined the specific vulnerabilities and attack techniques that emerge within LLM-powered AI agent communications. Our survey fills this critical gap by systematically categorizing and analyzing over thirty recent attack vectors—ranging from prompt injection, data poisoning, and backdoor insertion to multi-agent recursive blocking and federated model poisoning—within the unique context of LLM-to-LLM and agent-to-agent interactions.
Unlike prior works, which often address either generic LLM safety [wang2025comprehensive], tool integration [ferrag2025llm], or security in isolated modalities [yang2024mma, li2024seeing], our survey provides a vertical deep dive into vulnerabilities across multiple communication layers (e.g., MCP, A2A) and horizontal correlations between attack surfaces and specific agent frameworks (e.g., CrewAI, LangChain, AutoGen). Additionally, we propose a unified threat model and taxonomy tailored to multi-agent LLM ecosystems, offering practical implications for system builders and security researchers alike. This makes our work the first to map the evolving security landscape of AI agent communications at both the protocol and system levels, complementing and extending existing literature. In summary, while existing surveys provide valuable insights into communication protocols, safety taxonomies, and cybersecurity applications, none integrate these layers into a single, unified threat framework. Our survey builds on these foundations by correlating vulnerabilities across input, model, system, and protocol layers, thereby offering an end-to-end view of LLM-agent security that complements and extends previous literature.
@@
Perez et al.[perez2022ignore] 2022 Prompt‐Based Direct Prompt Injection (PromptInject) GPT‐3 (transformer‐based LLM) Discussion of need for stronger protective measures; no concrete defense tested Simple handcrafted inputs can induce goal hijacking and prompt leaking; low‐skill attackers can exploit stochastic behavior to misalign GPT‐3
@@
Pedro et al.[pedro2023prompt] 2023 Prompt‐Based Prompt‐to‐SQL (P2SQL) Injection Langchain‐based web apps integrating LLMs Four defense strategies integrated into Langchain (e.g., input sanitization, query validation) Seven SOTA LLMs are widely susceptible; defense strategies significantly reduce the success of SQL injection via prompts
@@
Bagdasaryan et al.[bagdasaryan2023abusing] 2023 Prompt‐Based Indirect & Compositional Prompt Injection Multimodal LLMs (LLaVA, PandaGPT) Recommendation for enhanced detection of adversarial perturbations in images/audio Adversarial perturbations embedded in images or sounds can steer multimodal LLM outputs without modifying textual prompt
@@
Zhan et al.[zhan2025adaptive] 2025 Prompt‐Based Adaptive Indirect Prompt Injection LLM agents using external tools Emphasis on designing defenses resilient to adaptive attacks; no novel defense, but evaluation methodology Eight existing defenses bypassed with adaptive strategies; attack success rate >50% for each defense
@@
Jiang et al.[jiang2025deceiving] 2025 Jailbreaking Compositional Instruction Attack (CIA) GPT‐4, GPT‐3.5, Llama2‐70B Intent‐Based Defense (IBD) to detect harmful intents CIA achieves >95% ASR; IBD reduces ASR by >74% by identifying underlying malicious intent
@@
Wei et al.[wei2023jailbreak] 2023 Jailbreaking In‐Context Demonstrations Attack (ICA) Various aligned LLMs In‐Context Defense (ICD) providing refusal examples Harmful demonstrations can increase jailbreak success; ICD effectively counters ICA by reducing success rates
@@
Shayegani et al.[shayegani2023jailbreak] 2023 Jailbreaking Cross‐Modality Jailbreak Attack Vision‐Language Models (e.g., CLIP‐integrated VLMs) Need for new alignment strategies in multi‐modal systems Adversarial images targeting toxic embeddings, combined with text, can break VLM alignment without model access
@@
Anil et al.[anil2024many] 2024 Jailbreaking Long‐Context Jailbreak Attack Closed‐weight LLMs (e.g., Anthropic, OpenAI) Recommendation for enhanced safety in long‐context models Attack success follows a power law as demonstration count increases; hundreds of shots suffice to manipulate behavior
@@
Liu et al.[liu2023autodan] 2023 Jailbreaking Automated Jailbreak Prompt Generation (AutoDAN) Aligned LLMs fine‐tuned with human feedback Hierarchical genetic algorithm framework AutoDAN produces stealthy, semantically meaningful prompts; outperforms prior methods in cross‐model transfer and stealth
@@
Yu et al.[yu2023gptfuzzer] 2023 Jailbreaking Jailbreak Fuzzing (GPTFuzz) ChatGPT, LLaMA‐2, Vicuna, etc. Judgment model to evaluate and refine templates >90% success rates across multiple LLMs; framework scales better than manual template crafting
@@
Schwartz et al.[schwartz2025graph] 2025 Jailbreaking Graph of Attacks with Pruning (GAP) GPT‐3.5 Automatic seed prompt generation; PromptGuard fine‐tuning for detection GAP achieves 92% ASR using 54% of queries required before; PromptGuard accuracy improves from 5.1% to 93.89% on Toxic‐Chat
@@
Dong et al.[dong2025fuzz] 2025 Jailbreaking JailFuzzer (LLM‐Agent‐Driven) Text‐to‐Image generative models LLM‐based guided mutation and oracle evaluation High success rates with minimal queries; generates natural, semantically coherent jailbreak prompts in black‐box settings
@@
Wang et al.[wang2023adversarial] 2023 Adversarial & Evasion adversarial In‐Context Learning (advICL) In‐Context Learning in LLMs Transferable‐advICL to craft demos for unseen inputs Robustness decreases as demonstration count increases; advICL and Transferable‐advICL mislead model on unseen inputs
@@
Zhuang et al.[zhuang2023pilot] 2023 Adversarial & Evasion Query‐Free Adversarial Attack Stable Diffusion (T2I model with CLIP encoder) Insights on text encoder robustness; suggests encoder hardening Five‐character perturbations can drastically alter generated images; targeted attacks steer specific content manipulations
@@
Yang et al.[yang2024mma] 2024 Adversarial & Evasion Multimodal Adversarial Attack (MMA‐Diffusion) T2I Models (open and commercial) Need for combined textual and visual defense strategies Bypasses prompt filters and post‐hoc safety checkers using text + visual manipulation; NSFW content generated despite safeguards
@@
Chen et al.[chen2025aeia] 2025 Adversarial & Evasion Active Environment Injection Attack (AEIA) MLLM‐Based Agents in Mobile OS AEIA‐MN framework to evaluate and reveal vulnerabilities Maximum 93% ASR on AndroidWorld when combining interface and reasoning gaps; advanced MLLMs remain highly vulnerable
@@
Patlan et al.[patlan2025real] 2025 Adversarial & Evasion Context Manipulation Attack Web3‐Based AI Agents Fine‐tuning methods more robust than prompt defenses CrAIBench finds prompt defenses ineffective; fine‐tuning reduces ASR significantly while preserving utility in single‐step tasks
@@
Input Manipulation Attacks Prompt-Based Attacks Direct Prompt Injection (Perez et al. [perez2022ignore]) Prompt-to-SQL Injection (Pedro et al. [pedro2023prompt]) Indirect & Compositional Prompt Injection (Bagdasaryan et al. [bagdasaryan2023abusing]) Adaptive Indirect Prompt Injection (Zhan et al. [zhan2025adaptive]) Toxic Agent Flow Attack [milanta2025githubmcp] Jailbreaking Attacks Compositional Instruction Attack (Jiang et al. [jiang2025deceiving]) In-Context Demonstrations (Wei et al. [wei2023jailbreak]) Long-Context Attacks (Anil et al. [anil2024many]) Automated Jailbreak Prompt Generation (Liu et al. [liu2023autodan]) Jailbreak-Fuzzing Attack (Dong et al. [dong2025fuzz]) Adversarial & Evasion Attacks Adv-ICL Attacks (Wang et al. [wang2023adversarial]) Query-Free Attack (Zhuang et al. [zhuang2023pilot]) Multimodal Adversarial Attack (Yang et al. [yang2024mma]) Active Environment Injection Attack (Chen et al. [chen2025aeia]) Context Manipulation Attack (Patlan et al. [patlan2025real])